You make a $9,000 deposit to an investment account today. The investment earns 7% p.a compounding monthly for the first 12 months, then it earns 10%p.a compounding monthly for the next 2 years. At the end of the 3 years the balance in the account is (to the nearest whole dollar; don't use $ sign or commas)
Answer #1

Let us find value of account at end of 1st year
FV = PV(1+r)^n
=9000(1+7%/12)^1*12
=9000(1+0.5833%)^12
=9000(1.005833)^12
=9000(1.072286)
=9650.572&
Now let us find value at end of year 3
FV = 9650.572(1+10%/12)^2*12
=9650.572(1+0.8333%)^24
=9650.572(1.008333)^24
=9650.572(1.220381)
=11777.38$